[figure: 9.7 cm]

DE HAAS–VAN ALPHEN EFFECT IN REGa_{3} COMPOUNDS

A. Czopnik^{a}, G.E. Grechnev^{b}, V.B. Pluzhnikov^{a,b,c} aW. Trzebiatowski Institute of Low Temperature and Structure Research, P.O.Box 1410, 50-950 Wrocław, Poland bB. Verkin Institute for Low Temperature Physics, Kharkov, 310164, Ukraine cInternational Laboratory of High Magnetic Fields and Low Temperatures, Gajowicka 95, 53-529 Wrocław, Poland

The electronic structure and cyclotron masses of TmGa_{3} , ErGa_{3} and LuGa_{3} compounds are evaluated by means of the de Haas–van Alphen effect study in strong enough magnetic fields, where magnetization is almost saturated. Also, the pressure effect on dHvA oscillations has been examined in ErGa_{3} and LuGa_{3} . Concurrently, ab initio calculations of the volume–dependent band structures have been carried out for these compounds, including ferromagnetic configuration phases of TmGa_{3} and ErGa_{3} , in an external magnetic field by employing relativistic versions of the LMTO method with the Zeeman term included. The mass enhancement factors λ were evaluated based on observed (m_{c}^{∗}) and calculated (m_{c}^{b}) cyclotron masses. The estimated magnetic contributions to λ in ErGa_{3} and TmGa_{3} appeared to be large and magnetic field dependent. A surprisingly large and intriguing pressure effect on cyclotron masses has been observed in ErGa_{3} . As a whole, the calculated pressure derivatives of the dHvA frequencies appeared to be in agreement with the experimental ones, though a nature of some discrepancies is not clear. A discussion is given on the role of different interactions (CEF, exchange splitting, magnetic–quadrupolar excitations, spin waves) in the effects revealed for REGa_{3}.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : A. Czopnik

Address for correspondence : W. Trzebiatowski Institute of Low Temperature and Structure Research, P.O.Box 1410, 50-950 Wrocław, Poland

Email address : czopnik@int.pan.wroc.pl